A=(a^2 ab ac Ab b^2 BC Ac BC c^2) then a^2+b^2+c^2=1 then A^2=?

Grade:12th pass

1 Answers

Vikas TU
5 years ago
 A2= a4+a2b2+a2c2+a3b+ab3+abc2+a3c+ab2c+ac3+a3b+ab3+abc2+a2b2+b4+b2c2+a2bc+b3c+a3c+ab2c+ac3+a2bc+b3c+bc3+a2c2+b2c2+c4
=a2(a2+b2+c2)+a3b+ab3+abc2+a3c+ab2c+ac3+a3b+ab3+abc2+b2(a2+b2+c2)+a2bc+b3c+a3c+ab2c+ac3+a2bc+b3c+bc3+c2(a2+b2+c2)
We know that a2+b2+c2=1 from the identitiy.
=a2+a3b+ab3+abc2+a3c+ab2c+ac3+a3b+ab3+abc2+b2+a2bc+b3c+a3c+ab2c+ac3+a2bc+b3c+bc3+c2
=1+a3b+ab3+abc2+a3c+ab2c+ac3+a3b+ab3+abc2+a2bc+b3c+a3c+ab2c+ac3+a2bc+b3c+bc3